Hi, I am not able to find a correct answer for the following question. Can someone please solve it? Thanks!

Hankins, Inc., is considering a project that will result in initial aftertax cash savings of $6.1 million at the end of the first year, and these savings will grow at a rate of 3 percent per year indefinitely. The firm has a target debt–equity ratio of .60, a cost of equity of 13 percent, and an aftertax cost of debt of 5.5 percent. The cost-saving proposal is somewhat riskier than the usual project the firm undertakes; management uses the subjective approach and applies an adjustment factor of +3 percent to the cost of capital for such risky projects.

WACC             %

What is the maximum cost the company would be willing to pay for this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
Present value            $
Answer #1

1.
=0.6/1.6*5.5%+1/1.6*13%=10.188%

2.
=6.1/(10.188%+3%-3%)=$59.8744 million or $59,874,361.99